# Foundation Models for Music: A Survey

### v1

@@ V-A6 Future Work
Advances in music modelling require more expansive and diverse datasets. Current resources, despite their extensive size, predominantly feature Western popular music, limiting model versatility. To develop robust models that grasp and generate music authentically across various cultures, it’s crucial to integrate a broad spectrum of global musical styles and genres into these datasets.
The current Multimodal Music Instruct dataset needs substantial enhancement to support foundation models effectively. It should expand not only in volume but also in the complexity of its content, particularly in areas of sequential MIR downstream tasks like beat recognition and chord progression. Besides, introducing a chain-of-thought component is essential for enabling models to perform complex reasoning and explain the musical theories underpinning their responses, thus fostering deeper analytical capabilities in music understanding.
@@
Last but not least, there is a need to find reasonable ratios of balancing downstream tasks and pre-training data to develop effective tuning strategies for injecting new knowledge or expertise while avoiding catastrophic forgetting, e.g., allowing the GPT to learn more about music, adapting the model to new music genres, or changes it more suitable to user preferences.
@@ V-B Evaluation
Music understanding and generation are of increasing interest to both academic and industry researchers and have the potential to broadly impact the economics and culture of music. However, evaluation remains an open and difficult research problem: models are not evaluated in the same downstream tasks, data Leakage, model bias, etc. In the following subsection, we will review commonly used evaluation protocols and metrics for both music understanding and music generation.
@@ V-B1 Evaluation for Music Understanding
We start by considering traditional MIR tasks. In this scenario, the goal is to measure music understanding by determining a model’s ability to correctly recognise specific musical properties at different levels of granularity, from frame- to track-level, by adopting the canonical MIR evaluation approach for the corresponding task. For models operating in the audio domain, tasks typically included are a subset of: key and mode detection [GDSB23a, LYZ+24], tempo detection [GDSB23a, LYZ+24], genre classification [GDSB23a, LYZ+24, DESW23], instrument recognition [GDSB23a, LYZ+24, DESW23], music tagging [LYZ+24], emotion recognition [LYZ+24], pitch detection [LYZ+24, DESW23], beat tracking [LYZ+24], and, in the less common performance domain, vocal technique detection [LYZ+24], singer identification [LYZ+24], and performance assessment [ZLD24, ZJJH21], among others. MIR-based evaluation can be executed in two distinct ways: via probing [LYZ+24, DESW23, CDL21] or via text-based instruction [GDSB23a]. Probing is used on models that produce audio representations: the audio backbone, which is either kept frozen or fine-tuned, is used as a feature extractor, and a lightweight probing network, typically a shallow MLP, is trained on top to solve the downstream task of interest. Instead, in models equipped with a text interface, MIR-based evaluation can be realised by prompting the model with an adequate text input (e.g. “What is the key of this song?”), and casting the output text (e.g “This song is in the key of F minor.”) to the corresponding label in the dataset.
Recently, there have been several efforts to unify this kind of evaluation methods across the full spectrum of music understanding tasks, by creating benchmarks that either include the music domain as a subset of interest [TSK+22] or fully specialise in it [YML+24]. The most notable of these is MARBLE [YML+24], which proposes a unified protocol for evaluating music audio representations across tasks at different levels of hierarchy, including high-level description (key detection, music tagging, genre classification, and emotion recognition), score-level (pitch tracking, beat tracking, melody extraction, chord estimation, lyrics transcription), performance-level (ornament and technique detection), and acoustic-level tasks (singer identification, instrument classification, source separation).
@@
In the context of multimodal foundation models, and beyond traditional MIR tasks, there is a similarly wide variety of tasks that are useful in evaluating music understanding. Among these, one that stands out as a common approach to evaluate fundamental capabilities in this domain is cross-modal retrieval, defined as the task of retrieving data samples of one modality (e.g. a music clip), based on a query sample of another modality (e.g. a textual description or a video). Models that are amenable to this type evaluation are most typically those trained to learn joint embedding spaces between two or more music-related data modalities, such as music-text contrastive models [HJL+22, MBQF22a, WYTS23a], or those trained to learn correspondences between music and images or videos [MSSR23, TRH23, SAFN24, SVRS22, VDG19, YZT+24]. In the case of music-text retrieval, this evaluation approach is not only useful for directly measuring performance on the task itself, but it also offers a framework for evaluating models on music classification tasks by casting them as text-based retrieval. In this case, cross-modal retrieval results in a further variation of the MIR-based evaluation approach presented earlier [DWCN23].
A further type of multimodal evaluation comprises tasks that are centered around language generation. As such, this is only suitable to music understanding models that take audio or (audio, text) pairs as inputs and generate text outputs. Under this setting, music understanding is measured by assessing language outputs that typically encompass several musical concepts. In practice, this is articulated in two subtasks: music captioning [TYS+24a, LHSS23a, DML+24, GDSB23a] and music question answering (MQA), which can be either open-ended [TYS+24a, HLSS23b, LHSS23a, DML+24] (sometimes also referred to as music reasoning [GDSB23a]), or multiple-choice [WMB+24]. In music captioning and open-ended MQA, the model is instructed to produce a language output that either describes the audio input (e.g. “Describe the contents of the provided audio in detail.”) or answers a question about it (e.g. “What are some possible uses for this music in a film or TV show?”). Performance is then measured in one of three ways: via automatic metrics, via human evaluation or via another LLM, using a strategy called LLM-as-a-judge [ZCS+23]. In the first case, automatic evaluation typically consists of computing text generation metrics such as BLEU [PRWZ02], METEOR [BL05], ROUGE [Lin04], and BERT-score [ZKW+20], which measure the linguistic or semantic similarity between the generated and reference text. We note that MQA evaluation bares similarities with the flavour of MIR-based evaluation that makes use of (question, answer) pairs. A key difference, however, is that the former is characterised by open-ended questions, while the latter focuses on a single musical aspect and assumes a one-to-one mapping between model output and ground truth.
@@
Despite the flourishing of research in this field, evaluating music understanding models in a consistent and reliable fashion remains a challenge due to limited benchmarks dedicated to the music domain. As a consequence, existing works follow inconsistent evaluation practices which make use of different datasets, metrics, experimental settings, and even task formulations. This becomes particularly prominent in language-based evaluation, where, as highlighted earlier in this section, there is little in the way of standardisation, and automatic metrics become unsuitable due to open-ended nature of the tasks [GDSB23a]. Text generation metrics are also known to deviate from human judgement when there isn’t a sufficient number of reference sentences or, similarly, when there are several potentially correct outputs, as is often the case in music understanding. This has prompted some to conduct human studies to obtain more reliable evaluations on open-ended tasks [GDSB23a].
Finally, comprehensive evaluation should go beyond assessing task-specific performance, and include other key factors like robustness, safety, efficiency, and alignment with human preferences. While this is becoming commonplace in other domains, it remains unexplored in the context of music understanding. .
@@ V-B2 Objective Evaluation for Music Generation
Similar to the field of music generation itself, research on objective evaluation for music generation models is in a nascent stage–The field has yet to agree on a set of standard metrics and/or benchmarks, and new metrics have been constantly proposed in an ad-hoc manner with new methods. Objective evaluation of machine-generated music is challenging for several main reasons: (i) The aesthetics of music is multifaceted, e.g., harmony, melody, instrumentation, and how they are assembled together, which are both hard to quantify individually, and highly interrelated with one another. (ii) Music generation is, in fact, a family of tasks that can take a variety of user inputs, ranging from a short text description [ADB+23, CKG+24] to fine-grained musical concepts like melody [CKG+24], chords [vRBKH23], beats [WDWB24], and even music composition with rendering (performance) objective [PCCKW23, ZCCC+24], hence necessitating task-specific metrics.
@@
A functional type (or type in short) defines the high-level workflow to map a set of generated (and/or reference) samples to a metric value. Let one sample of reference (i.e., human-composed) music be denote by $\bm{x}$, and one sample of generated music by $\bm{\tilde{x}}$. Also, let $X=\{\bm{x}^{(i)}\}_{i=1}^{N}$ denote the set of $N$ reference samples, and $\tilde{X}=\{\bm{\tilde{x}}^{(i)}\}_{i=1}^{\tilde{N}}$ the set of $\tilde{N}$ generated samples. We further define $f(\bm{x})$ or $f(\bm{\tilde{x}})$ to be any function that extracts an attribute or some features from music, either global ones (e.g., a text description for the music) or local ones (e.g., a chord sequence), which may sometimes also be independently given by users as conditions, and $\mathrm{score}(\cdot,\cdot)$, or $\mathrm{score}(\cdot)$ to be a deterministic function that computes the reported metric value. Commonly adopted metrics may be categorised as follows:
@@
The purpose of a metric are associated with either of the two following major focuses:
@@
  • Control-focused: Such metrics examine how well a generative model adheres to the conditions given by users. Frequently-used conditions are text descriptions, musical genres, emotions, instrumentation, tempos, chord sequences, melodies, etc.
@@
As for metrics that focus on one specific aspect, the Structureness indicator (type: S-I) [WY20] is a representative example, which examines the presence of repetitive structures across all time granularities. Structureness indicator leverages Fitness Scape-plots [MJ12] as $f(\bm{x})$, and then applies a $\max$ operation on the scape-plot as the $\mathrm{score(\cdot)}$ function, which is meant to describe how much the most-repeated excerpt within a specified time granularity (e.g., 10$\sim$20 seconds) is repeated throughout the entire audio.
@@
  • CLAP score (CLAP, type: P-I, $\uparrow$) [WCZ+23b] depends on a text encoder and an audio encoder as feature extractors $f_{1},f_{2}$. The two encoders should have been jointly pre-trained via contrastive learning [vdOLV18] to form an aligned latent feature space. Then, CLAP score is simply the cosine similarity between the encoded features of the input text, and those of the generated audio. MuLan score [HJL+22] is an almost identical metric, only with a different bi-encoder backbone.
@@
To evaluate adherence to more fine-grained and musically-centered types of controls, the metrics are often proposed in individual works and hence not yet standardised. Commonly considered musical controls and their corresponding metrics are as follows:
@@
  • Groove consistency (type: S-I) [WY20] divides each bar into segments (e.g., in 32th note increments), and construct a binary groove vector for each bar by checking whether each segment has note onsets. Then, some similarity (e.g., Hamming similarity) between groove vectors of neighboring (or all) bar pairs is computed.
@@
  • Similarity error (type: P-A, $\downarrow$): was proposed by [YLW+22] and encompasses both pitch and rhythm to examine whether generations are structurally similar to human-composed music. It first constructs a note set for each bar, whose elements describe the (pitch, duration, onset time relative to downbeat) of each note, and then computes the mean intersection-over-union (IoU) similarity of all bar pairs that are $t$ bars apart (where $t$ is preset). Finally, the difference between the mean IoUs between human-composed and generated pieces are taken.
@@ V-B3 Subjective Evaluation for Music Generation
Due to the highly artistic and subjective nature of music, subjective studies, i.e., listening tests with human participants, still remain the gold standard of evaluation music generation models [YL20, JYL23]. Regardless of whether the model operates in audio or symbolic domain, the samples presented to participants are typically (synthesised) audio waveforms. Below we explain two common protocols for subjective studies.
@@
On the other hand, compared to pair-wise comparisons, more fine-grained aspects are often considered. Although various terms and questions have been used, the fine-grained aspects may roughly be group into several categories:
@@ VII Conclusion
Our exploration of FMs’ representation in music has revealed that current models predominantly focus on a limited spectrum of musical representations such as spectrum, waveform, MIDI, and ABC notation. Besides, we observe there exists a considerable gap in fully exploiting multimodal representations that unify symbolic music, audio, text, and music score images within FM development.
@@
Moreover, the survey highlights several significant technical advances and remaining challenges in leveraging LLMs and LDMs for music, including pre-training paradigms, domain adaptation, in-context learning, model architectures, controllability, music agents and scaling law. Training paradigms, including contrastive learning, generative pretraining, and mask language modelling, have shown potential in learning from large-scale, unsupervised music data. However, all of these models require fine-tuning on downstream tasks, and models developed with the combination of pre-training and instruction tuning are yet to be fully realised. Besides, combining the music domain to develop pre-training targets has also not been fully discovered.
Domain adaptation techniques such as adapter and prefix tuning have proven effective for fine-tuning pre-trained language models on downstream music tasks, enhancing the multimodal capabilities of these models without extensive retraining. These methods facilitate the integration of music-specific knowledge into LLMs, allowing for nuanced understanding and generation of musical content. However, the research field still lacks a comprehensive understanding of the potential of MLP, Q-former or other architecture on music FMs.
@@
The architectural of most current foundation models, primarily based on the Transformer architecture, supports extensive scalability but faces challenges with long-sequence music data. Innovations in model architecture are required to better handle the complexities of musical structure and longer context windows. Although the GPT-4 and Claude-3 can handle more than 100k tokens, there is still a lot of room for improvement in the ability to handle long sequences. Specific architectural adjustments or algorithms may be needed to enhance the modelling and exploitation of long contextual information. Another issue is that current foundation models are mainly decoder-only architectures, and training such models places high demands on the quality of the music-text tokeniser. Whether the Encoder-decoder architecture is an alternative is also worth exploring.
@@
Lastly, music agents represent a promising area of development, transitioning from semi-autonomous to fully autonomous systems that can interact with users, manage complex workflows, and integrate various musical tasks and modalities seamlessly. These agents are also poised to revolutionise how we interact with and create music, offering personalized and adaptive music experiences.

### v2

@@ V-A6 Future Work
Advances in music modelling require more expansive and diverse datasets. Current resources, in spite of their extensive size, predominantly feature Western popular music, limiting model versatility. To develop robust models that grasp and generate music authentically across various cultures, it’s crucial to integrate a broad spectrum of global musical styles and genres into these datasets.
The current Multimodal Music Instruct dataset needs substantial enhancement to support foundation models effectively. It should expand not only in volume but also in the complexity of its content, particularly in areas of sequential MIR downstream tasks like beat recognition and chord progression. Furthermore, introducing a chain-of-thought component is essential for enabling models to perform complex reasoning and explain the musical theories underpinning their responses, thus fostering deeper analytical capabilities in music understanding.
@@
Last but not least, there is a need to find reasonable ratios of balancing downstream tasks and pre-training data to develop effective tuning strategies for injecting new knowledge or expertise while avoiding catastrophic forgetting, e.g., allowing the GPT to learn more about music, adapting the model to new music genres, or make it more suitable to user preferences.
@@ V-B Evaluation
Music understanding and generation are of increasing interest to both academic and industry researchers and have the potential to broadly impact the economics and culture of music. However, evaluation remains an open and difficult research problem, facing challenges like inconsistent downstream evaluation conditions, data leakage, model bias, etc. In the following subsection, we will review commonly used evaluation protocols and metrics for both music understanding and music generation.
@@ V-B1 Evaluation for Music Understanding
We start by considering traditional MIR tasks. In this scenario, the goal is to measure music understanding by determining a model’s ability to correctly recognise specific musical properties at different levels of granularity, from frame to track level, by adopting the canonical MIR evaluation approach for the corresponding task. For models operating in the audio domain, tasks typically included are a subset of key and mode detection [GDSB23a, LYZ+24], tempo detection [GDSB23a, LYZ+24], genre classification [GDSB23a, LYZ+24, DESW23], instrument recognition [GDSB23a, LYZ+24, DESW23], music tagging [LYZ+24], emotion recognition [LYZ+24], pitch detection [LYZ+24, DESW23], beat tracking [LYZ+24], and, in the less common performance domain, vocal technique detection [LYZ+24], singer identification [LYZ+24], and performance assessment [ZLD24, ZJJH21], among others. MIR-based evaluation can be executed in two distinct ways: via probing [LYZ+24, DESW23, CDL21] or via text-based instruction [GDSB23a]. Probing is used on models that produce audio representations: the audio backbone, which is either kept frozen or fine-tuned, is used as a feature extractor, and a lightweight probing network, typically a shallow MLP, is trained on top to solve the downstream task of interest. Instead, in models equipped with a text interface, MIR-based evaluation can be realised by prompting the model with an adequate text input (e.g. “What is the key of this song?”), and casting the output text (e.g “This song is in the key of F minor.”) to the corresponding label in the dataset.
Recently, there have been several efforts to unify this kind of evaluation methods across the full spectrum of music understanding tasks, by creating benchmarks that either include the music domain as a subset of interest [TSK+22] or fully specialise in it [YML+24, PAJB23]. The most extensive of these is MARBLE [YML+24], which proposes a unified protocol for evaluating music audio representations across tasks at different levels of hierarchy, including high-level description (key detection, music tagging, genre classification, and emotion recognition), score-level (pitch tracking, beat tracking, melody extraction, chord estimation, lyrics transcription), performance-level (ornament and technique detection), and acoustic-level tasks (singer identification, instrument classification, source separation).
@@
In the context of multimodal foundation models and beyond traditional MIR tasks, there is a similarly wide variety of tasks that are useful in evaluating music understanding. Among these, one that stands out as a common approach to evaluate fundamental capabilities in this domain is cross-modal retrieval, defined as the task of retrieving data samples of one modality (e.g. a music clip), based on a query sample of another modality (e.g. a textual description or a video). Models that are amenable to this type evaluation are most typically those trained to learn joint embedding spaces between two or more music-related data modalities, such as music-text contrastive models [HJL+22, MBQF22a, WYTS23a], or those trained to learn correspondences between music and images or videos [MSSR23, TRH23, SAFN24, SVRS22, VDG19, YZT+24]. In the case of music-text retrieval, this evaluation approach is not only useful for directly measuring performance on the task itself, but it also offers a framework for evaluating models on music classification tasks by casting them as text-based retrieval. In this case, cross-modal retrieval results in a further variation of the MIR-based evaluation approach presented earlier [DWCN23].
A further type of multimodal evaluation comprises tasks that are centred around language generation. As such, this is only suitable to music understanding models that take audio or (audio, text) pairs as inputs and generate text outputs. Under this setting, music understanding is measured by assessing language outputs that typically encompass several musical concepts. In practice, this is articulated in two subtasks: music captioning [TYS+24a, LHSS23a, DML+24, GDSB23a] and music question answering (MQA), which can be either open-ended [TYS+24a, HLSS23b, LHSS23a, DML+24] (sometimes also referred to as music reasoning [GDSB23a]), or multiple-choice [WMB+24]. In music captioning and open-ended MQA, the model is instructed to produce a language output that either describes the audio input (e.g. “Describe the contents of the provided audio in detail.”) or answers a question about it (e.g. “What are some possible uses for this music in a film or TV show?”). Performance is then measured in one of three ways: via automatic metrics, via human evaluation or via another LLM, using a strategy called LLM-as-a-judge [ZCS+23]. In the first case, automatic evaluation typically consists of computing text generation metrics such as BLEU [PRWZ02], METEOR [BL05], ROUGE [Lin04], and BERT-score [ZKW+20], which measure the linguistic or semantic similarity between the generated and reference text. We note that MQA evaluation bears similarities with the flavour of MIR-based evaluation that makes use of (question, answer) pairs. A key difference, however, is that the former is characterised by open-ended questions, while the latter focuses on a single musical aspect and assumes a one-to-one mapping between model output and ground truth.
@@
Despite the flourishing of research in this field, evaluating music understanding models in a consistent and reliable fashion remains a challenge due to limited benchmarks dedicated to the music domain. As a consequence, existing works follow inconsistent evaluation practices which make use of different datasets, metrics, experimental settings, and even task formulations. This becomes particularly prominent in language-based evaluation, where, as highlighted earlier in this section, there is little in the way of standardisation, and automatic metrics become unsuitable due to the open-ended nature of the tasks [GDSB23a]. Text generation metrics are also known to deviate from human judgement when there isn’t a sufficient number of reference sentences or, similarly, when there are several potentially correct outputs, as is often the case in music understanding. This has prompted some to conduct human studies to obtain more reliable evaluations on open-ended tasks [GDSB23a].
Finally, comprehensive evaluation should go beyond assessing task-specific performance and include other key factors like robustness, safety, efficiency, and alignment with human preferences. While this is becoming commonplace in other domains, it remains unexplored in the context of music understanding.
@@ V-B2 Objective Evaluation for Music Generation
Similar to the field of music generation itself, research on objective evaluation for music generation models is in a nascent stage–The field has yet to agree on a set of standard metrics and/or benchmarks, and new metrics have been constantly proposed in an ad-hoc manner with new methods. Objective evaluation of machine-generated music is challenging for several main reasons: (i) The aesthetics of music is multifaceted, e.g., harmony, melody, instrumentation, and how they are assembled together, which are both hard to quantify individually and highly interrelated with one another. (ii) Music generation is, in fact, a family of tasks that can take a variety of user inputs, ranging from a short text description [ADB+23, CKG+24] to fine-grained musical concepts like melody [CKG+24], chords [vRBKH23], beats [WDWB24], and even music composition with rendering (performance) objective [PCCKW23, ZCCC+24], hence necessitating task-specific metrics.
@@
A functional type (or type in short) defines the high-level workflow to map a set of generated (and/or reference) samples to a metric value. Let one sample of reference (i.e., human-composed) music be denoted by $\bm{x}$, and one sample of generated music by $\bm{\tilde{x}}$. Also, let $X=\{\bm{x}^{(i)}\}_{i=1}^{N}$ denote the set of $N$ reference samples, and $\tilde{X}=\{\bm{\tilde{x}}^{(i)}\}_{i=1}^{\tilde{N}}$ the set of $\tilde{N}$ generated samples. We further define $f(\bm{x})$ or $f(\bm{\tilde{x}})$ to be any function that extracts an attribute or some features from music, either global ones (e.g., a text description for the music) or local ones (e.g., a chord sequence), which may sometimes also be independently given by users as conditions, and $\mathrm{score}(\cdot,\cdot)$, or $\mathrm{score}(\cdot)$ to be a deterministic function that computes the reported metric value. Commonly adopted metrics may be categorised as follows:
@@
The purpose of a metric is associated with either of the two following major focuses:
@@
  • Control-focused: Such metrics examine how well a generative model adheres to the conditions given by users. Frequently used conditions are text descriptions, musical genres, emotions, instrumentation, tempos, chord sequences, melodies, etc.
@@
As for metrics that focus on one specific aspect, the Structureness indicator (type: S-I) [WY20] is a representative example, which examines the presence of repetitive structures across all time granularities. The Structureness indicator leverages Fitness Scape-plots [MJ12] as $f(\bm{x})$, and then applies a $\max$ operation on the scape-plot as the $\mathrm{score(\cdot)}$ function, which is meant to describe how much the most-repeated excerpt within a specified time granularity (e.g., 10$\sim$20 seconds) is repeated throughout the entire audio.
@@
  • CLAP score (CLAP, type: P-I, $\uparrow$) [WCZ+23b] depends on a text encoder and an audio encoder as feature extractors $f_{1},f_{2}$. The two encoders should have been jointly pre-trained via contrastive learning [vdOLV18] to form an aligned latent feature space. Then, CLAP score is simply the cosine similarity between the encoded features of the input text and those of the generated audio. MuLan score [HJL+22] is an almost identical metric, only with a different bi-encoder backbone.
@@
To evaluate adherence to more fine-grained and musically-centred types of controls, the metrics are often proposed in individual works and hence not yet standardised. Commonly considered musical controls and their corresponding metrics are as follows:
@@
  • Groove consistency (type: S-I) [WY20] divides each bar into segments (e.g., in 32th note increments), and construct a binary groove vector for each bar by checking whether each segment has note onsets. Then, some similarity (e.g., Hamming similarity) between groove vectors of neighbouring (or all) bar pairs is computed.
@@
  • Similarity error (type: P-A, $\downarrow$): was proposed by [YLW+22] and encompasses both pitch and rhythm to examine whether generations are structurally similar to human-composed music. It first constructs a note set for each bar, whose elements describe the (pitch, duration, onset time relative to downbeat) of each note, and then computes the mean intersection-over-union (IoU) similarity of all bar pairs that are $t$ bars apart (where $t$ is preset). Finally, the difference between the mean IoUs between human-composed and generated pieces is taken.
@@ V-B3 Subjective Evaluation for Music Generation
Due to the highly artistic and subjective nature of music, subjective studies, i.e., listening tests with human participants, still remain the gold standard of evaluation music generation models [YL20, JYL23]. Regardless of whether the model operates in the audio or symbolic domain, the samples presented to participants are typically (synthesised) audio waveforms. Below we explain two common protocols for subjective studies.
@@
On the other hand, compared to pair-wise comparisons, more fine-grained aspects are often considered. Although various terms and questions have been used, the fine-grained aspects may roughly be grouped into several categories:
@@ VII Conclusion
Our exploration of FMs’ representation in music has revealed that current models predominantly focus on a limited spectrum of musical representations such as spectrum, waveform, MIDI, and ABC notation. Besides, we observe a considerable gap in fully exploiting multimodal representations that unify symbolic music, audio, text, and music score images within FM development.
@@
Moreover, the survey highlights several significant technical advances and remaining challenges in leveraging LLMs and LDMs for music, including pre-training paradigms, domain adaptation, in-context learning, model architectures, controllability, music agents and scaling law. Training paradigms, including contrastive learning, generative pretraining, and mask language modelling, have shown potential in learning from large-scale, unsupervised music data. However, all of these models require fine-tuning on downstream tasks, and models developed with the combination of pre-training and instruction tuning are yet to be fully realised. Besides, combining the music domain to develop pre-training targets has not been fully discovered.
Domain adaptation techniques such as adaptor and prefix tuning have proven effective for fine-tuning pre-trained language models on downstream music tasks, enhancing the multimodal capabilities of these models without extensive retraining. These methods facilitate the integration of music-specific knowledge into LLMs, allowing for nuanced understanding and generation of musical content. However, the research field still lacks a comprehensive understanding of the potential of MLP, Q-former or other architecture on music FMs.
@@
The architectural of most current foundation models, primarily based on the Transformer architecture, supports extensive scalability but faces challenges with long-sequence music data. Innovations in model architecture are required to better handle the complexities of musical structure and longer context windows. Although the GPT-4 and Claude-3 can handle more than 100k tokens, there is still much room for improvement in the ability to handle long sequences. Specific architectural adjustments or algorithms may be needed to enhance the modelling and exploitation of long contextual information. Another issue is that current foundation models are mainly decoder-only architectures, and training such models places high demands on the quality of the music-text tokeniser. Whether the Encoder-decoder architecture is an alternative is also worth exploring.
@@
Lastly, music agents represent a promising area of development, transitioning from semi-autonomous to fully autonomous systems that can interact with users, manage complex workflows, and integrate various musical tasks and modalities seamlessly. These agents are also poised to revolutionise how we interact with and create music, offering personalised and adaptive music experiences.